# TimeDART: A Diffusion Autoregressive Transformer for Self-Supervised Time Series Representation

### v3

@@ 1 Introduction
Figure 1: Two-stage training process. Pre-training: TimeDART teaches an encoder to learn temporal representations through a noising process and autoregressive denoising. Fine-tuning: The pre-trained encoder is adapted with task-specific heads for downstream applications.
@@
The analysis of time-series data has become increasingly critical across diverse application domains, including healthcare (Cheng et al., 2024b), finance (Black & Scholes, 1973) and energy management (Zhou et al., 2024). These scenarios often involve massive amounts of unlabeled real-world data. To analyze and leverage such data, self-supervised learning has emerged as a widely adopted approach (Zhang et al., 2024b). This method extracts valuable knowledge from abundant unlabeled data, which can be subsequently transferred to boost downstream task performance (Park et al., 2024; Dong et al., 2024b).
@@
Reviewing previous studies (Zhang et al., 2024b), existing methods primarily fall into three categories. The first category is masked autoencoders, which focus on reconstructing masked or corrupted parts of the input and excel at learning the underlying patterns of the time series (He et al., 2022; Zerveas et al., 2021). In addition, Cheng et al. (2023b) proposed TimeMAE to mitigate inconsistencies between pre-training and fine-tuning caused by newly injected masked embeddings. Contrastive-based methods, as the second category, specialize in sequence-level modeling and distinguish between similar and dissimilar time-series segments (Eldele et al., 2021; Tonekaboni et al., 2021). Nonetheless, their emphasis on sequence-level modeling may limit their capacity to capture fine-grained temporal variations (Dong et al., 2024a). The third category is autoregressive methods, which are well-suited for modeling left-to-right sequential relationships and have promising scale-up potential. (Liu et al., 2024b). However, autoregressive methods tend to overfit noise and anomalies in time series (Lim & Zohren, 2021), thereby exacerbating error accumulation, which may help explain the limited emphasis on such methods in self-supervised representation learning.
We argue that an ideal self-supervised approach should address the aforementioned challenges, including simultaneously capturing long-term dynamic evolution and subtle local patterns in a unified manner. While autoregressive-based self-supervised learning aligns well with the natural left-to-right trend of time series, it is prone to overfitting the common noise and anomalies in the data. To solve these problems, we explicitly introduce a diffusion and denoising process in the auto-regressive optimization, enhancing data diversity and increasing pre-training difficulty. This design forces the model to learn useful information from noise during autoregressive pre-training, while also avoiding the problem of diffusion models overly emphasizing local structures at the expense of global dependencies.
@@
Building upon the analysis above, we propose TimeDART, a novel self-supervised time series representation framework that unifies two powerful generative paradigms to learn more transferable representations, as illustrated in Figure 1. Specifically, we initialize a vanilla Transformer encoder and partition the input time-series data into non-overlapping patches. To prevent information leakage during autoregressive optimization, we apply a causal mask within the encoder. Instead of directly optimizing the differences between the encoder output and original patches, we independently add noise to each patch and use a cross-attention-based denoising network, which leverages preceding sequence information to guide the denoising. This design enables TimeDART to effectively capture both global and local sequence features in a coherent way, achieving superior performance across nine publicly available datasets on downstream forecasting and classification tasks.
@@ 2 Related Work
2.1 Self-supervised Learning in Time Series.
Self-supervised learning has become a powerful pre-training approach in fields like natural language processing and computer vision. Unlike supervised learning, which relies on labeled data, self-supervised methods generate supervision from the data’s structure, often through pretext tasks. In time series, self-supervised learning faces unique challenges due to the sequential and temporal nature of the data. Current approaches (Zhang et al., 2024b) can be broadly categorized into three main paradigms: masked autoencoders, contrastive learning and autoregressive methods.
Masked Autoencoders
As a fundamental approach in self-supervised representation learning, mask autoencoders typically involve reconstructing masked or corrupted inputs, encouraging the model to learn meaningful representations. Masked time series modeling, introduced by TST (Zerveas et al., 2021), predicts missing time points from available data. Methods like STEP (Shao et al., 2022), PatchTST (Nie et al., 2022), and CrossTimeNet (Cheng et al., 2024a) extend this approach by operating on sub-series, reducing computational costs while capturing local information. More recent work, such as TimeMAE (Cheng et al., 2023b), introduces decoupled masked autoencoders to address inconsistencies between the pre-training and fine-tuning phases. Additionally, SimMTM (Dong et al., 2024b) improves masked time-series modeling by recovering missing time points through weighted aggregation of neighbors, leveraging the manifold structure of the data.
Contrastive Learning
@@
Autoregressive Methods
In time series, autoregressive methods originated from ARIMA (Box et al., 2015), which combines autoregressive (AR) and moving average (MA) models with differencing. Later, with the rise of RNNs, THOC (Shen et al., 2020) introduced a self-supervised pretext task for multi-resolution single-step forecasting called Temporal Self-Supervision. Recently, inspired by pre-trained large language models (Brown et al., 2020), the potential of autoregressive models in time series has gained attention, leveraging their generalization ability and task versatility (Liu et al., 2024b). These language models have even been adapted as autoregressive predictors (Liu et al., 2024a) to achieve arbitrary input-output mappings. Nevertheless, the full potential of autoregressive methods in time series representation learning and various downstream tasks remains to be discovered. To address this gap, TimeDART uses autoregressive optimization during pretraining, improving time series representations and enabling strong performance across downstream tasks.
2.2 Diffusion Models in Time Series.
In recent years, Denoising Diffusion Probabilistic Models (Ho et al., 2020; Li et al., 2024) have become powerful tools for time series modeling (Lin et al., 2023; Meijer & Chen, 2024) due to their unique advantages in fine-grained temporal modeling. Early models, such as CSDI (Tashiro et al., 2021), avoid autoregressive inference while incorporating additional input masking. TimeGrad (Rasul et al., 2021) introduced autoregressive denoising with Langevin sampling, enhancing multivariate prediction. More recently, conditional diffusion models like TimeDiff (Shen & Kwok, 2023) have improved time series prediction by leveraging external information to guide the denoising process. However, these methods primarily focus on probabilistic time series prediction tasks, aiming to generate high-quality time series data, but their direct application to downstream tasks like time series forecasting often yields suboptimal results (Zhang et al., 2024a). TimeDART uniquely integrates the denoising diffusion process within a self-supervised framework, enabling inner-patch modeling that preserves the subtle local patterns time series data.
@@ 3.1 The Proposed TimeDART
TimeDART involves three modules: normalization and patch embedding, causal Transformer encoder and and a patch-level diffusion denoising process.
@@ 3.1.1 Normalization and Embedding
We use patches as the basic modeling unit, capturing richer local information for more comprehensive representations (Cheng et al., 2023a). Applying noise and denoising to individual points could lead to excessive sensitivity, whereas patches offer more stable embeddings. To preserve the autoregressive property and prevent information leakage, we set the patch length $P$ equal to the stride $S$, ensuring non-overlapping patches that respect the autoregressive assumption. For simplicity, we assume $L$, the time-series length, is divisible by $P$, resulting in $N=\frac{L}{P}$ patches, which reduces computational complexity and facilitates the processing of longer sequences.
Each patch (referred to as a clean patch) is then passed through a linear embedding layer, transforming it into a high-dimensional representation:
@@ 3.1.2 Causal Transformer Encoder
We initialize a vanilla Transformer encoder as the representation network. During pre-training, we prepend a learnable start-of-sequence (SOS) embedding to the clean patch representations while exclude the final one. To incorporate positional information, we apply sinusoidal positional encoding after the embedding layer. Following this, we use a causal mask $M$ in the processing layer, limiting the visibility of each patch to itself and the prior patches. Consequently, the causal encoder network can be expressed as:
$\bm{z}^{in}_{1:N}$ $=\operatorname{Concat}[SOS,~{}\bm{z}_{1:N-1}]+\operatorname{PE}_{1:N},$ (2)
$f(\bm{z}^{in}_{1:N})$ $=\operatorname{Encoder}(\bm{z}^{in}_{1:N},~{}M),$ (3)
@@ 3.1.3 Patch-Level Diffusion and Denoising
$\hat{\bm{z}}^{in}_{1:N}=\operatorname{Embedding}(\hat{\bm{x}}_{1:N})+\operatorname{PE}_{1:N},$ (6)
@@
The reverse process is handled by the denoising decoder which takes the encoder output as keys and values, while the noise-added patch embeddings serve as queries. A self-only mask is applied to the decoder to ensure that the $j$-th input in the noise-added sequence attends only to the $j$-th encoder output. Informed by the causal mask, the encoder output at position $j$ aggregates information from clean patches at positions $1$ to $j-1$, enabling autoregressive optimization. Finally, deep representations are mapped back to the original space via linear projection. The reverse process is expressed as:
$z_{j}^{out}=g(\hat{z}^{in}_{j},~{}f(\bm{z}^{in}_{1:j-1})),\quad 1\leq j\leq N,$ (7)
where $g(\cdot)$ denotes the processing of the encoder output and noise-added patch embeddings by the denoising decoder.
@@ 3.2 Self-supervised Optimization Objective
Our self-supervised optimization objective minimizes the diffusion loss, equivalent to the Evidence Lower Bound (ELBO (Ho et al., 2020)). The final loss is:
$\mathcal{L}_{diff}=\sum_{j=1}^{N}\mathbb{E}_{\epsilon,q(x_{j}^{0})}\left[||x_{j}^{0}-g(\hat{z}^{in}_{j},~{}f(\bm{z}^{in}_{1:j-1}))||^{2}\right].$ (8)
We replace conventional MSE loss with associated diffusion loss, enabling the model to better express a multimodal belief over time-series data (Li et al., 2024). This formulation is embedded in the autoregressive optimization objective, which effectively captures both long-term dynamic evolution and subtle local patterns. The detailed derivation can be found in Appendix C.
@@ 4 Experiments
We perform extensive experiments on various datasets across two main time series tasks, namely forecasting and classification, to demonstrate the effectiveness of TimeDART. Furthermore, we conduct ablation studies that emphasize the capacity of the autoregressive mechanism, while validating the necessity of the diffusion and denoising process within this. We also analyze the impact of key parameters on the architecture, demonstrating its robustness.
@@ 4.1 Experimental Setup
Table 2: Multivariate time series forecasting results. All results are averaged Mean Squared Error (MSE) and Mean Absolute Error (MAE) from 4 different predicted windows of $\{12,24,36,48\}$ for PEMS datasets and $\{96,192,336,720\}$ for others. The best results are in bold and the second best are underlined. Full results are detailed in Appendix D.
@@
In TimeDART, a simple linear transformation is used for the embedding layer. Additionally, to ensure the autoregressive mechanism functions correctly, it is crucial that patches remain non-overlapping to prevent information leakage. Therefore, we prioritize using a vanilla non-overlapping patch embedding layer and a Transformer encoder, rather than relying on existing backbones such as PatchTST (Nie et al., 2022) or iTransformer (Liu et al., 2023). To investigate whether TimeDART maintains consistent performance across different backbones, we conducted additional experiments by replacing the Transformer encoder with a causal TCN under specific configurations, as it also satisfies the aforementioned requirements. Regardless of the backbone, we use a vanilla Transformer decoder as the denoising decoder, with parameters identical to those of the Transformer encoder, except for the number of layers.
@@
We compare TimeDART against four state-of-the-art self-supervised baseline methods, including a contrastive learning method: CoST (Woo et al., 2022), and three masked modeling methods: SimMTM (Dong et al., 2024b), TimeMAE (Cheng et al., 2023b) and the self-supervised version of PatchTST (Nie et al., 2022). Since we use a vanilla Transformer encoder as the backbone instead of what is proposed in existing supervised methods, we believe it is essential to compare with advanced supervised methods that feature carefully designed architectures. For the forecasting task, we compare with the supervised version of PatchTST (Nie et al., 2022) and the linear-based method DLinear (Zeng et al., 2023). For the classification task, we compare with FormerTime (Cheng et al., 2023a), which employs a hierarchical Transformer network as the backbone. More implementation details can be found in Appendix B.1.
@@
To ensure the fairness of the experiments, we use a unified embedding layer and Transformer encoder as the representation network for all self-supervised methods, keeping model-specific parameters consistent. In particular, for forecasting experiments, we apply the channel-independent configuration (Nie et al., 2022) for rubustness (Han et al., 2024) and experiments under cross-domain setting. For supervised methods, we respect their original implementations without modifying model-specific parameters, making adjustments only to some non-core training parameters for alignment.
In addition, for the forecasting task, we adopt two experimental settings: in-domain and cross-domain. In the in-domain setting, we follow the common look-back window length of $L=336$ and predicted windows $H\in\{96,192,336,720\}$ for all forecasting datasets except the PEMS dataset, where the corresponding $L=96$ and $H\in\{12,24,36,48\}$. On the other hand, in the cross-domain setting, we mix all forecasting datasets and perform general pre-training on a consistent look-back window of $L=336$, while keeping the same setup as the aforementioned setting during downstream fine-tuning. To emphasize the improvements achieved through pre-training, we also conducted experiments with Random Init., which means randomly initializing the representation network and directly fine-tuning without pre-training. More details on the fairness of the experiments can be found in Appendix B.2.
Table 4: Multivariate time series classification results. Results are are reported as accuracy (Acc.) and F1 score (F1). The best results are in bold and the second best are underlined.

### v5

@@ 1 Introduction
Figure 1: Two-stage pre-training and fine-tuning process. a): TimeDART teaches an encoder to learn temporal representations through a noising process and autoregressive denoising. b): The pre-trained encoder is adapted with task-specific heads for downstream applications.
@@
The analysis of time-series data has become increasingly critical across diverse application domains, including healthcare (Cheng et al., 2024), finance (Black & Scholes, 1973) and energy management (Zhou et al., 2024). These scenarios often involve massive amounts of unlabeled real-world data. To analyze and leverage such data, self-supervised learning has emerged as a widely adopted approach (Zhang et al., 2024b). This approach extracts valuable knowledge from abundant unlabeled data, which can be subsequently transferred to boost downstream task performance (Park et al., 2024; Dong et al., 2024b).
@@
Reviewing previous studies (Zhang et al., 2024b), existing methods for time series self-supervised learning primarily fall into three categories. The first category is masked autoencoders, which focus on reconstructing masked or corrupted parts of the input and excel at learning the underlying patterns of the time series (He et al., 2022; Zerveas et al., 2021). A common challenge with these models is the inconsistency between pre-training and fine-tuning due to newly injected masked embeddings (Cheng et al., 2023b). Contrastive-based methods, as the second category, specialize in sequence-level modeling and distinguish between similar and dissimilar time-series segments (Eldele et al., 2021; Tonekaboni et al., 2021). Nonetheless, their emphasis on sequence-level modeling may limit their capacity to capture fine-grained temporal variations (Dong et al., 2024a). The third is autoregressive methods, which are well-suited for modeling left-to-right sequential relationships and have promising scale-up potential (Liu et al., 2024c). However, these methods tend to overfit noise and anomalies in time series data (Lim & Zohren, 2021). Moreover, Using the common squared error loss in autoregressive time series modeling tends to make the current value’s prediction collapse into a Gaussian distribution centered on past values with a fixed variance (Li et al., 2024b), which deviates significantly from real-world scenarios.
We argue that an ideal self-supervised approach should address the above challenges, that is, how to simultaneously capture long-term dynamic evolution and subtle local patterns in a unified manner. Although the autoregressive self-supervised learning shows significant potential due to its natural alignment with the left-to-right temporal dynamics, a key challenge lies in its limited capacity for modeling fine-grained patterns and its tendency to collapse into a Gaussian distribution. We attempt to address this by incorporating a diffusion denoising process within the autoregressive optimization framework. This design teaches the model to learn useful representations from the explicit added noise during autoregressive pre-training, and rationalizing the squared error of this regression problem using the diffusion loss.
@@
Building upon the analysis above, we propose TimeDART, a novel self-supervised time series representation framework that unifies two powerful generative paradigms, as illustrated in Figure 1. Specifically, we initialize a vanilla Transformer encoder and partition the input time-series data into non-overlapping patches. To prevent information leakage during autoregressive optimization, we apply a causal mask within the encoder. We independently add noise to each patch and use a cross-attention-based denoising decoder, which leverages preceding sequence information to guide the denoising process. TimeDART achieves superior performance across nine publicly available datasets on downstream forecasting and classification tasks.
@@ 2 Related Work
2.1 Self-supervised Learning in Time Series
Self-supervised learning has emerged a powerful pre-training approach in fields like natural language processing and computer vision (Brown et al., 2020; He et al., 2022). Unlike supervised learning, which relies on labeled data, self-supervised methods generate supervision from the data’s structure. Applying this to time-series data presents unique challenges due to its sequential and temporal characteristics. Current approaches (Zhang et al., 2024b; Cheng et al., 2025a) can be categorized into three main paradigms: masked reconstruction, contrastive discrimination and autoregressive prediction.
Masked Reconstruction
As a fundamental approach in self-supervised representation learning, mask reconstruction typically involve reconstructing masked or corrupted inputs, encouraging the model to learn meaningful representations. Masked time series modeling, introduced by TST (Zerveas et al., 2021), predicts missing time points from available data. Methods like STEP (Shao et al., 2022), PatchTST (Nie et al., 2023), and CrossTimeNet (Cheng et al., 2025b) extend this approach by operating on sub-series, reducing computational costs while capturing local information. More recent work, such as TimeMAE (Cheng et al., 2023b), introduces decoupled masked autoencoders to address inconsistencies between the pre-training and fine-tuning phases. Additionally, SimMTM (Dong et al., 2024b) improves masked time-series modeling by recovering missing time points through weighted aggregation of neighbors, leveraging the manifold structure of the data.
Contrastive Discrimination
@@ 2.1 Self-supervised Learning in Time Series
Autoregressive Prediction
In time series, autoregressive prediction originated from ARIMA (Box et al., 2015), which combines autoregressive (AR) and moving average (MA) models with differencing. Later, with the rise of RNNs, THOC (Shen et al., 2020) introduced a self-supervised pretext task for multi-resolution single-step forecasting called Temporal Self-Supervision. Recently, inspired by pre-trained large language models (Brown et al., 2020), the potential of autoregressive models in time series has gained attention, leveraging their generalization ability and task versatility (Liu et al., 2024c). These language models have even been adapted as autoregressive predictors (Liu et al., 2024b) to achieve arbitrary input-output mappings. Nevertheless, the full potential of autoregressive prediction in time series representation learning and various downstream tasks remains to be discovered. To address this gap, TimeDART uses autoregressive optimization during pretraining, improving time series representations and enabling strong performance across downstream tasks.
2.2 Diffusion Models in Time Series
In recent years, Denoising Diffusion Probabilistic Models (Ho et al., 2020; Li et al., 2024b) have become powerful tools for time series modeling (Lin et al., 2024; Meijer & Chen, 2024) due to their unique advantages in fine-grained temporal modeling. Early models, such as CSDI (Tashiro et al., 2021), avoid autoregressive inference while incorporating additional input masking. TimeGrad (Rasul et al., 2021) introduced autoregressive denoising with Langevin sampling, enhancing multivariate prediction. More recently, conditional diffusion models like TimeDiff (Shen & Kwok, 2023) have improved time series prediction by leveraging external information to guide the denoising process. However, these methods primarily focus on probabilistic time series prediction tasks, aiming to generate high-quality time series data, but their direct application to downstream tasks like time series forecasting often yields suboptimal results (Zhang et al., 2024a). TimeDART uniquely integrates the denoising diffusion process within a self-supervised framework, enabling inner-patch modeling that preserves the subtle local patterns time series data.
@@ 3.1 The Proposed TimeDART
TimeDART involves three modules: normalization and patch embedding, causal Transformer encoder and a patch-level diffusion denoising process.
@@ 3.1.1 Normalization and Embedding
We use patches as our fundamental modeling unit to capture richer local information and generate more comprehensive representations (Cheng et al., 2023a). This approach offers more stable embeddings compared to applying noise and denoising to individual data points, which can lead to excessive sensitivity. To maintain the autoregressive property and prevent information leakage, we ensure non-overlapping patches by setting the patch length $P$ equal to the stride $S$. For simplicity, we assume the time-series length $L$ is divisible by $P$, yielding $N=\frac{L}{P}$ patches. This reduces computational complexity and streamlines the processing of longer sequences. Subsequently, each clean patch is transformed into a high-dimensional representation via a linear embedding layer:
@@ 3.1.2 Causal Transformer Encoder
We initialize a vanilla Transformer encoder (Vaswani et al., 2017) as the representation network. During pre-training, we prepend a learnable start-of-sequence (SOS) embedding to the clean patch representations while exclude the final one. To incorporate positional information, we apply sinusoidal positional encoding after the embedding layer. Following this, we use a causal mask $M$ in the processing layer, limiting the visibility of each patch to itself and the prior patches. Finally, the causal encoder network can be expressed as:
$\bm{z}^{\text{in}}_{1:N}$ $=\operatorname{Concat}[SOS,~{}\bm{z}_{1:N-1}]+\operatorname{PE}_{1:N},$ (2)
$f(\bm{z}^{\text{in}}_{1:N})$ $=\operatorname{Encoder}(\bm{z}^{\text{in}}_{1:N},~{}M),$ (3)
@@ 3.1.3 Patch-Level Diffusion and Denoising
$\hat{\bm{z}}^{\text{in}}_{1:N}=\operatorname{Embedding}(\hat{\bm{x}}_{1:N})+\operatorname{PE}_{1:N},$ (6)
@@
The reverse process is handled by the denoising decoder which takes the encoder output as keys and values, while the noise-added patch embeddings serve as queries. A self-only mask is applied to the decoder to ensure that the $j$-th input in the noise-added sequence attends only to the $j$-th encoder output. Informed by the causal mask, the encoder output at position $j$ aggregates information from clean patches at positions $1$ to $j-1$, enabling autoregressive optimization. The reverse process is expressed as:
$z_{j}^{\text{out}}=g(\hat{z}^{\text{in}}_{j},~{}f(\bm{z}^{\text{in}}_{1:j-1})),\quad 1\leq j\leq N,$ (7)
where $g(\cdot)$ denotes the processing of the encoder output and noise-added patch embeddings by the denoising decoder. Finally, deep representations are mapped back to the original space via linear projection:
$x_{j}^{\text{out}}=\operatorname{Projector}(z_{j}^{\text{out}}),$ (8)
where $\bm{x}^{\text{out}}_{1:L}\in\mathbb{R}^{C\times L}$, with the same shape as the input $\bm{x}_{1:L}$.
@@ 3.2 Self-supervised Optimization Objective
Traditional time series autoregressive pre-training objective is typically a simple mean squared error (MSE), aiming to minimize the discrepancy between the regressed and true patches. This can be formalized as follows:
$\mathcal{L}_{\text{mse}}\propto\sum_{j=1}^{N}||x_{j}^{0}-\operatorname{Projector}(f(\bm{z}^{in}_{1:j-1}))||^{2}.$ (9)
While straightforward, this MSE objective fundamentally assumes that the underlying data distribution is Gaussian. This assumption is evident when viewing MSE through the lens of maximum likelihood estimation:
$\frac{1}{2\sigma^{2}}||x_{j}^{0}-\operatorname{Projector}(f(\bm{z}^{in}_{1:j-1}))||^{2}$ $=$ (10)
$-\log\mathcal{N}(x_{j}^{0};\operatorname{Projector}(f(\bm{z}^{in}_{1:j-1}$ $)),\sigma^{2})+C,$
where $\sigma$ is a constant and $C$ is a constant determined by $\sigma$. If it is valid, sampling $x_{j}^{0}$ could be achieved via $x_{j}^{0}=\operatorname{Projector}(f(\bm{z}^{in}_{1:j-1}))+\sigma\epsilon$, where $\epsilon$ is standard Gaussian noise. This implies time series observations are either very noisy or follow a unimodal, bell-shaped distribution. Real-world time series, however, often display complex, multimodal behaviors and intricate dependencies that a simple Gaussian assumption poorly captures.
To overcome these limitations, TimeDART replaces MSE with a diffusion loss, allowing the model to express a richer, multimodal belief over time series data. This self-supervised diffusion loss, formally equivalent to the Evidence Lower Bound (Ho et al., 2020; Chen et al., 2024):
$\mathcal{L}_{\text{diff}}=\sum_{j=1}^{N}\mathbb{E}_{\epsilon,q(x_{j}^{0})}\left[||x_{j}^{0}-g(\hat{z}^{in}_{j},~{}f(\bm{z}^{in}_{1:j-1}))||^{2}\right].$ (11)
This diffusion loss, embedded in the autoregressive optimization objective, effectively captures both long-term dynamic evolution and subtle local patterns. The core idea is to make the square error a reasonable loss by gradually adding and then removing noise (Su, 2024; Li et al., 2024a). Refer to Appendix C for the detailed derivation.
@@ 4 Experiments
We demonstrate TimeDART’s effectiveness in forecasting and classification across diverse datasets. Ablation studies confirm the power of its autoregressive mechanism and the critical role of diffusion denoising process, further showing its robustness through parameter analysis.
@@ 4.1 Experimental Setup
Table 2: Multivariate time series forecasting results. All results are averaged MSE and Mean Absolute Error (MAE) from 4 different predicted windows of $\{12,24,36,48\}$ for PEMS datasets and $\{96,192,336,720\}$ for others. The best results are in bold and the second best are underlined. Full results are detailed in Appendix D.
@@
In TimeDART, a simple linear transformation is used for the embedding layer. Additionally, to ensure the autoregressive mechanism functions correctly, it is crucial that patches remain non-overlapping to prevent information leakage. Therefore, we prioritize using a vanilla non-overlapping patch embedding layer and a Transformer encoder, rather than relying on existing backbones such as PatchTST (Nie et al., 2023) or iTransformer (Liu et al., 2024a). To investigate whether TimeDART maintains consistent performance across different backbones, we conducted additional experiments by replacing the Transformer encoder with a causal TCN under specific configurations, as it also satisfies the aforementioned requirements. Regardless of the backbone, we use a vanilla Transformer decoder as the denoising decoder, with parameters identical to those of the Transformer encoder, except for the number of layers.
@@
We compare TimeDART against four state-of-the-art self-supervised baseline methods, including a contrastive learning method: CoST (Woo et al., 2022), and three masked modeling methods: SimMTM (Dong et al., 2024b), TimeMAE (Cheng et al., 2023b) and the self-supervised version of PatchTST (Nie et al., 2023). Since we use a vanilla Transformer encoder as the backbone instead of what is proposed in existing supervised methods, we believe it is essential to compare with advanced supervised methods that feature carefully designed architectures. For the forecasting task, we compare with the supervised version of PatchTST (Nie et al., 2023) and the linear-based method DLinear (Zeng et al., 2023). For the classification task, we compare with FormerTime (Cheng et al., 2023a), which employs a hierarchical Transformer network as the backbone. More implementation details can be found in Appendix B.1.
@@
To ensure experimental fairness, we employ a unified embedding layer and Transformer encoder with consistent model-specific parameters. Specifically, for forecasting experiments, we implement a channel-independent configuration (Nie et al., 2023) to enhance robustness (Han et al., 2024) and facilitate cross-domain analysis. In addition, we conduct forecasting experiments under two settings: in-domain and cross-domain. In the in-domain setup, a look-back window of $L=336$ and prediction horizons $H\in\{96,192,336,720\}$ are used for all datasets, except PEMS where $L=96$ and $H\in\{12,24,36,48\}$. For the cross-domain setting, all forecasting datasets are mixed for general pre-training with a consistent look-back window of $L=336$, while downstream fine-tuning maintains as in-domain settings. To highlight benefits of pre-training, we also include baseline Random Init., involving direct fine-tuning without pre-training. Supervised methods retain their original implementations, with only non-core training parameters adjusted for alignment. Further details on experimental fairness are provided in Appendix B.2.
Table 4: Multivariate time series classification results. Results are are reported as Accuracy (Acc.) and Macro-F1 (F1). The best results are in bold and the second best are underlined.